# Score-based Neural Ordinary Differential Equations for Computing Mean Field Control Problems

### v1

@@ B.2 Pseudo-code for algorithms
MFC problem equation 28 equation 24, $N_{t}$, $N_{z}$, learning rate, weight parameter $\lambda$, number of iterations
@@
              Compute $\mathcal{L}_{\text{HJB}t_{j}}$ through equation 86
@@
         compute $z_{t_{j+1}}^{(n)}$, $\tilde{l}_{t_{j+1}}^{(n)}$, $s_{t_{j+1}}^{(n)}$, $H_{t_{j+1}}^{(n)}$ through the forward Euler scheme equation 9, equation 11, equation 12, and equation 13
@@
MFC problem equation 80 equation 81, $N_{t}$, $N_{z}$, learning rate, weight parameter $\lambda$, number of iterations
@@
              Compute $\mathcal{L}_{\text{HJB}t_{j}}$ through equation 93
@@
         compute $z_{t_{j+1}}^{(n)}$ $s_{t_{j+1}}^{(n)}$, $H_{t_{j+1}}^{(n)}$ through the forward Euler scheme equation 9, equation 12, and equation 13
@@
MFC problem equation 19 equation 20, $N_{t}$, $N_{z}$, network structure equation 8, learning rate, number of iterations, partition of the total time interval $\{I_{m}\}_{m=1}^{N_{\text{stage}}}$
@@ B.3 The errors in the numerical examples
The errors we present in the tables in section 5 and Appendix C are the average errors over $10$ independent runs. Note that there are two sources that contribute to these three errors $\text{err}_{\rho}$, $\text{err}_{f}$, and $\text{err}_{s}$. We take the terminal time $t_{t_{N}}=t_{\text{end}}$ as an example. The first source is the discretization error for the state dynamic, characterized by $z_{t_{\text{end}}}-z_{t_{N_{t}}}$, where $z_{t_{\text{end}}}$ is the end point of the continuous dynamic equation 3a and $z_{t_{N_{t}}}$ is its time discretization. The second source comes from the training error of the neural network $f(\cdot,\cdot;\theta)$. When a function is steep, especially the score function $\nabla\log\rho$ in a non-Gaussian example, a small discretization error could be magnified by a large factor due to this large condition number. This is the reason why the score errors for the double well potential function example in Appendix C.4 are larger compared with other results.
The errors for the double well potential example. For the double well potential example, it is hard to obtain a reference solution for $t\in(0,t_{\text{end}})$ (cf. Appendix C.4). As an alternative, we compute the error of the velocity field at initial and terminal time $t=0,t_{\text{end}}$ through
@@
where $A(\cdot)$ and $B(\cdot)$ denote the true solutions. Here, we do not present the error of $\theta^{C}$ for two reasons. First, according to equation 84 and equation 91, the velocity field does not depend on $\theta^{C}$. Second, we observe from equation 86 and equation 93 that a constant shift of $\theta^{C}$ does not affect the regularization loss. Looking more carefully at equation 85, a constant shift of all the even (or odd) indices of $\theta^{C}$ also does not change the loss equation 86 or equation 93. Therefore, it is not meaningful to compute the error for $\theta^{C}$.
@@ B.4 Hyperparameters for the numerical examples
We set the diffusion coefficient $\gamma=1$ for all problems, except for the double well potential example, where $\gamma=0.1$. The reason for this difference is because if we set $\gamma=1$ in the double well example, the terminal distribution $\rho(1,\cdot)$ will be very flat and we will not have a clear bifurcation pattern. In the LQ example in Appendix C.3, we set $\beta=0.1$.
@@
For the neural network parametrization equation 8, we use $\tanh$ as the activation function. The widths of the network are $20$, $50$, and $200$ in $1$, $2$, and $10$ dimensions for the RWPO and LQ problems. For the double moon flow matching problem (in $2$ dimensions), the width is $200$. For the double well example, the widths are $50$, $100$, and $200$ in $1$, $2$, and $10$ dimensions. The double moon flow matching problem and the double well problem are harder due to the bifurcation pattern.
@@
Appendix C Additional numerical results
C.1 RWPO
For the RWPO problem, after the transformation to probability flow through equation 22, the problem becomes
subject to the transport equation
$\partial_{t}\rho(t,x)+\nabla_{x}\cdot(\rho(t,x)f(t,x))=0\,.$
The optimal density evolution is given by
and the optimal velocity field is
$f(t,x)=-\dfrac{x}{2(2-t)}\,.$
In addition to the errors and Figure 1, we present additional numerical results in Figure 2. The first, second, and third rows of Figure 2 show numerical results in $1$, $2$, and $10$ dimensions, respectively. The first column shows the curves of the cost function through training. The light blue shadows plot the standard deviation observed during $10$ independent test runs. The red dashed lines represent the cost under the optimal control field. We observe that our algorithm nearly reaches the optimal cost.
The fourth column in the first row of Figure 2 plots the score function (in blue) at $t=1$ using the data points $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$. This curve coincides with the true score function in orange. Note that the score function at $t=0$ is given, so we only plot the score at a terminal time. The second and third figures in the fourth column of Figure 2 show a density plot of the first dimension of the velocity field. These are probability density functions of $f(0,z_{0};\theta)$ and $f(0,z_{0})$, where $z_{0}\sim\rho_{0}$. The density functions are further approximated by the histogram of samples. Such techniques for visualization of high dimensional functions have been applied by Han et al. (2020). We observe that our neural network accurately captures the velocity fields.
C.2 Flow matching
In this section, we first present additional numerical results for the flow matching of the OU process. Then, we describe the details of the numerical implementation, such as the double moon potential, and present the numerical results.
Flow matching for OU process. The numerical results of flow matching for OU process are shown in Figure 3 and 4. Figure 3 shows the density evolution of the process in $1$ dimension, which is similar to the second plot in Figure 1. Our algorithm captures both the change of mean value and the shrink of variance accurately. Figure 4 shows the particle dynamic under the trained probability flow in $2$ dimensions and its comparison with the stochastic OU process. We add level sets of the density function with center $\mu(t)$ and radius $\frac{1}{4}\sigma(t)$, $\frac{1}{2}\sigma(t)$, and $\sigma(t)$ for better visualization, where $\sigma(t):=\det(\Sigma(t))^{\frac{1}{2d}}$ denotes the standard deviation. We pick points that are initialized within the circle with center $\mu(0)$ and radius $\sigma(0)$, and record the trajectories of these particles. According to Figure 4, our algorithm accurately captures the change of mean and variance of the OU process. Additionally, compared with the OU process, our probability flow ODE demonstrates structured behaviors.
Figure 4: Flow matching for OU process in $2$ dimension. First line: particle evolution under the probability flow $z_{t}$ with trained velocity field. Second line: simulation of the stochastic OU process. The circles in the figure are centered at $\mu(t)$, with radius $\frac{1}{4}\sigma(t)$, $\frac{1}{2}\sigma(t)$, and $\sigma(t)$. Our algorithm accurately captures the density evolution of the state dynamic, including the change of mean and variance. The probability flow demonstrate more structured behavior compared with the stochastic OU process.
Double moon example. In the example with double moon potential, given by
This example has been computed by Wang & Li (2022); Tan et al. (2023). We aim to learn the probability flow ODE of the overdamped Langevin dynamic
Next, we present the detailed implementation of the multi-stage splicing method for our double moon example. In this toy example, we partition the interval into two stages (sub-intervals) $[0,0.2]$ and $[0.2,0.4]$. We apply Algorithm 1 within the first interval. After training, we save the trained network as a warm-up (initialization) for the training in the next stage. We also record the particles $z_{t}^{(n)}$ at terminal time $t=0.2$ in the first stage, which serves as the initial distribution for training in the next stage.
In the second stage $t\in[0.2,0.4]$, the training process is similar. We inherit the network parameter $\theta$ from the previous stage as initialization. Also, the state particles $z_{t}^{(n)}$ at terminal time from the last stage are utilized as initialization distribution for the new stage. Note that starting form the second stage, there is no longer resampling because we only have finite samples at $t=0.2$. As a consequence, we may encounter the problem of overfitting, which is also known as model collapse. In this work, we apply a $L^{2}$ regularization with weight $0.1$ to avoid this issue. If we only have finite samples for the initial distribution, then regularization should also be added in the first stage of training. We summarize this multi-stage splicing method in Algorithm 4.
C.3 An LQ example with an entropy potential cost
We consider an LQ example in this section. This example was also studied by Lin et al. (2021). We minimize the cost functional
$\inf_{v}\int_{0}^{t_{\text{end}}}\int_{\mathbb{R}^{d}}\left(\frac{1}{2}\left|v(t,x)\right|^{2}+\frac{1}{2}|x|^{2}+\beta\log(\rho(t,x))\right)\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}^{d}}G(x)\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$
subject to
With a score substitution, the problem is equivalent to
$+\int_{\mathbb{R}^{d}}G(x)\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$
subject to
The numerical results are also presented in Figure 6. The first, second, and third rows show the results in $1$, $2$, and $10$ dimensions, respectively. The first column is the training curve for the cost functional, which is similar to the first column in Figure 2. Our algorithm gives the correct objective. The second column visualize the density function at terminal time $t_{\text{end}}$ using the data points $\{z^{(n)}_{t_{N_{t}}},\tilde{l}^{(n)}_{t_{N_{t}}}\}_{n=1}^{N_{z}}$. In the $1$ dimension (first row), we plot the density and compare it with true densities directly. In $2$ dimensions, we interpolate the data points and plot the density function $x_{1}\mapsto\rho(t_{\text{end}},(x_{1},0))$ in the figure. In $10$ dimensions, an interpolation is very hard to obtain, so we plot the density in term of the norm of $x$. Our algorithm captures the density function accurately. The third column visualize the score function at terminal time $t_{\text{end}}$ using the data points $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$. Again, we plot the score function directly in $1$ dimension. In $2$ and $10$ dimensions, we present a density plot for the first dimension of the score, which gives the probability density function of $\nabla_{z_{1}}\log\rho(t_{\text{end}},z_{t_{\text{end}}})$ (similar to the density plot in Appendix C.1). We observe that our algorithm captures the true density and score functions accurately.
errors $\text{err}_{\rho}$ $\text{err}_{f}$ $\text{err}_{s}$ cost gap
[…]
Table 5: Errors for LQ problem.
C.4 Double well potential
In this section we present an MFC example where the terminal cost is a potential function with the double well shape. The formulation is similar to the WPO, with the terminal cost given by
$G(x)=c\left|x-c_{1}\right|^{2}\left|x-c_{2}\right|^{2}\,,$ (100)
where $c\in\mathbb{R}^{+}$ and $c_{1},c_{2}\in\mathbb{R}^{d}$ are the minimizers of $G$. This function $G$ does not depend on the density $\rho$. The initial distribution is still a standard Gaussian distribution, i.e., $z_{0}\sim N(0,I_{d})$. As mentioned before, such problem will demonstrate a bifurcation phenomenon, which is hard to capture. We compute this example in $1$, $2$, and $10$ dimensions, and compute all errors in $1$ and $2$ dimensions. We take $c=\frac{1}{4}$ in equation 100 and set two centers as an all $-1$ vector for $c_{1}$ and an all $1$ vector for $c_{2}$.
Computing the reference solution. Unlike the example in section 5.1, we do not have an explicit solution. But fortunately, we are able to obtain a reference solution using the kernel formula proposed by Li et al. (2023) (equation (10), (11), and (14)). According to their derivations, the optimal density function is given by
$\rho(t,x)=\left(4\pi\gamma\frac{t(t_{\text{end}}-t)}{t_{\text{end}}}\right)^{-\frac{d}{2}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{\exp\left[-\frac{1}{2\gamma}\left(G(z)+\frac{|x-z|^{2}}{2(t_{\text{end}}-t)}+\frac{|x-y|^{2}}{2t}\right)\right]}{\int_{\mathbb{R}^{d}}\exp\left[-\frac{1}{2\gamma}\left(G(\tilde{y})+\frac{|y-\tilde{y}|^{2}}{2t_{\text{end}}}\right)\right]\,\mathrm{d}\tilde{y}}\rho(0,y)\,\mathrm{d}y\,\mathrm{d}z\,.$ (101)
Specifically, the density function as terminal time $t_{\text{end}}$ is
In addition, the solution to the classic HJB equation (cf. equation 75) is given by
with a terminal condition $\phi(T,x)=-G(x)$. With these expressions, we are able to obtain the score function via $\nabla_{x}\log\rho(t,x)=\nabla_{x}\rho(t,x)/\rho(t,x)$. Then, by Corollary 1 and equation 76, the optimal velocity is
$f(t,x)=\nabla_{x}\phi(t,x)-\gamma\nabla_{x}\log\rho(t,x)\,.$ (104)
However, when $t\in(0,t_{\text{end}})$, the numerical implementation for $\rho(t,x)$ and $\nabla_{x}\rho(t,x)$ through equation 101 involves three nested integrations in $\mathbb{R}^{d}$, which could potentially result in large errors. Therefore, we instead consider the errors at $t=0$ and $t_{\text{end}}$, where all expressions are relatively easy. At $t=0$, by equation 104, we have
At $t=t_{\text{end}}$, we can compute the score function through
Also, the terminal velocity is given by
Next, we apply the Riemann sum to approximate the expressions equation 105, equation 106, and equation 107 to obtain a reference solution in $1$ and $2$ dimensions. First, we define the following function
which has appeared multiple times. We approximated $h(y)$ numerically from the Riemann sum. In $1$ dimension, we use the trapezoid rule with step size $\Delta z=0.01$ to approximate the integrations and truncate the integration within $z\in[-6,6]$ when we compute the reference solution. In $2$ dimensions, similarly, we use a box with size $\Delta z_{1}\times\Delta z_{2}=0.01\times 0.01$ for Riemann sum and truncate the integration within the box $z\in[-6,6]\times[-6,6]$. We compute the value of $h(y)$ on all grid points within $[-4,4]$ in $1$ dimension and $[-4,4]\times[-4,4]$ in $2$ dimensions, and store the values. In this way, integrations w.r.t. $z$ in equation 105 and equation 106 are simulated numerically. Next, we further apply Riemann sum for integrations w.r.t. $y$ with the same step size, truncated for $y\in[-4,4]\,\text{or}\,[-4,4]\times[-4,4]$. In this way, we obtain the function for all $x$ on the grid points. The two bounds $6$ and $4$, and the step size $0.01$, are chosen such that the Riemann sums provide reasonable reference solutions for the problem. Finally, we make an interpolation in the dimensions $1$ or $2$, giving us the desired reference solution.
[…]
Figure 7: double well 1d. Left: true density and particle histogram at $T$; middle: density evolution compared with true density; right: score function at $T$.
[…]
The results are also shown in Figure 8 and 9. Our learned network successfully captures the velocity field $f$ and the score function which is shown in Figure 8. The first and second rows show each dimension of the velocity field at $t=0$ and $t=t_{\text{end}}=1$ and compare them with the true values. The third row shows the score function. The approximated score function is plotted using the data $\{(z^{(n)}_{t_{N_{t}}},s^{(n)}_{t_{N_{t}}})\}_{n=1}^{N_{z}}$, where we made a non-uniform $2d$ interpolation using the build-in function in Scipy. The reference solution is obtained through differentiation of equations (10) (11) and (14) given by Li et al. (2023) and then apply Riemann sum approximation. We observe that our algorithm correctly captures the velocity field and the score function.
The left plot in Figure 9 shows the scattered points at terminal time $t=1$, which concentrate at the two wells of the potential. We also add the level sets of $\rho(1,\cdot)$ at $\{0.01,0.05,0.1,0.2,0.4,0.6\}$ for better visualization. Note that the expression for $\rho(1,\cdot)$ is given by equation 102 instead of $\frac{1}{Z}\exp(-G(x)/\gamma)$. The plot in the middle shows the particle trajectories under the trained velocity field, which shows a bifurcation phenomenon.
Results in $10$ dimensions.
Finally, we also test our algorithm in $10$ dimensions. We do not have a reference solution in this case, but we observe that the particles go into two piles, as is shown in the right plot of Figure 9. In this plot, we project the $10$ dimensional particles into $2$ dimensions for visualization.
Appendix D remarks
@@ C.4 Double well potential
The score dynamic equation 3c. Boffi & Vanden-Eijnden (2023) proposed a similar equation (in equation (13)) before. But we have concrete numerical example. Also, we believe that we are the first to propose the ODE dynamic for the second-order score function equation 3d. This equation significantly enhances the accuracy of the regularized MFC solver, as demonstrated in Table 3 and 4.
Parametrization of neural network. In this paper, we parametrize $f$ as a neural network with a single (hidden) layer. As a consequence, the whole map from $z_{0}$ to $z_{t_{\text{end}}}$ is slightly different from the typical structure of a deep neural network, as described by Chen et al. (2018). The layer structure for $f$ is affine-activation-affine, so the composition of multiple layers becomes affine-activation-affine-affine-activation-affine. In contrast, a typical structure for a deep neural network is affine-activation-affine-activation-affine. The composition of two affine functions is still an affine function, so it is not a big difference.
Related work on theoretical MFC. From a theoretical perspective, convergence properties of MFC problems have been extensively studied by Carmona & Laurière (2021; 2022). In addition, Lacker (2017) investigated the limit behavior as the number of agents approaches infinity. The maximum principle, a crucial concept in optimal control, has also been extended to the mean field setting by Meyer-Brandis et al. (2012), offering an understanding of the theoretical underpinnings of MFC problems. The mean field game (MFG) problem (Cardaliaguet, 2010) is closely related to MFC. Cardaliaguet et al. (2019) generalize the HJB equation to the master equation, whose properties have been further investigated by Gangbo & Mészáros (2022); Gangbo et al. (2022). Bayraktar et al. (2024) study the convergence of particle system on Wasserstein space.
HJB regularization for MFC solver. For the HJB regularization technique in B.1, we used absolute value instead of a squared residual in equation 87 because it performs slightly better. In fact, both forms of regularization could significantly reduce errors.
As for the reason why it reduces errors, we recall that we minimize the discretized loss function in the original Algorithm 1. As a consequence, there is a trend of overfitting to the discretized loss. Adding a regularizer could resolve this overfitting issue. Such issues for discretizations are also discussed in Zhou & Lu (2024).
Computational cost for the score function via equation 17 and equation 18. When we compute the score function through equation 17, the term $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)$ could be tricky. Note that $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)=\nabla_{x}\det(\nabla_{x}T(t,x))/\det(\nabla_{x}T(t,x))$, and the $i$-th component of $\nabla_{x}\det(\nabla_{x}T(t,x))$ can be computed through
@@
Therefore, the $i$-th component of $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)$ in equation 17 is
@@
As a consequence, the cost for computing one single score function $s_{t}$ through equation 17 is either from computing the inverse of Jacobi matrix or solving the related linear system for each dimension. Classical numerical linear algebra methods suggest that the complexity is $\mathcal{O}(d^{3})$, and the cost for one score trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$.
We can also compute the score function through equation 18, where we assume that the width of the neural network is $\mathcal{O}(d)$, i.e., proportional the dimension. In this way, $\dfrac{\partial z_{t}}{\partial z_{0}}$ and $\dfrac{\partial l_{t}}{\partial z_{0}}$ are computed from auto-differentiations of the deep neural network functions in equation 15. The computational cost for a single score $s_{t_{j}}$ is $\mathcal{O}(j\,d^{3})$, because we need to compute the back propagation of the deep neural network equation 15 with depth $j$ using chain rule. However, the total cost for computing a whole trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$ rather than $\mathcal{O}(N_{t}^{2}\,d^{3})$ because we can efficiently store intermediate computations. For example, according to equation 9, we can compute
@@
with additional $\mathcal{O}(d^{2})$ operations if $\dfrac{\partial z_{t_{j}}}{\partial z_{0}}$ is stored. Similarly, by equation 10, we can compute
@@
with additional $\mathcal{O}(d^{2})$ operations provided that $\dfrac{\partial l_{t_{j}}}{\partial z_{0}}$ and $\dfrac{\partial z_{t_{j}}}{\partial z_{0}}$ are stored. Note that computing the inverse of the Jacobian matrix or solving the related linear systems requires $\mathcal{O}(d^{3})$ operations following classical numerical linear algebra methods. Therefore, computing $s_{t_{j+1}}$ through equation 18 requires additional $\mathcal{O}(d^{3})$ operations if we smartly record the results at $t_{j}$. Consequently, the total cost for computing a whole trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$. We also remark that Huang et al. (2024) give a derivation similar to equation 18.
Other methods to compute the score function. For the KDE method, an approximation for the density function is obtained through taking the kernel convolution with samples directly. Then, one compute its spatial derivative and obtain the score function via $\nabla_{x}\log\rho(t,x)=\nabla_{x}\rho(t,x)/\rho(t,x)$. However, the choice of kernel and bandwidth could be tricky, and the numerical error could be large, especially in high dimensions (Węglarczyk, 2018). Score matching (Hyvärinen & Dayan, 2005) is another commonly used method. With a small trick of integration by part and dropping a constant term that includes the integral with squared score functions, one only requires samples to train a network for the score function. However, training a neural network could be expensive, compared with the KDE method, which evaluates the score function directly.
Regularity for the MFC problem. Formula equation 3 requires third order derivatives for the velocity field $f$. As a consequence, we need regularity assumption for the MFC problem to guarantee that the ODE system is meaningful. In this work, we assume that $L$ and $F$ are $C^{1,2,2}$ Hölder continuous, and $G$ is $C^{2,2}$ Hölder continuous. According to Schauder estimates (Ladyzhenskaia et al., 1968, Chapter 4), these regularity assumptions guarantee that the velocity field has up to third-order derivatives, ensuring that the ODE system equation 3 is meaningful.
The multi-stage splicing method. For our multi-stage splicing method (cf. Appendix C.2), we partition the total time span evenly and pass the trained neural network from the previous stage into the next stage. When the dynamic becomes complicated, the length of the time interval could also be adapted to suit the problem (Zhou et al., 2021).
Also, the neural network from the previous stage is probably not an accurate velocity for the new stage, but it does contain more information than a random initialization. Therefore, this nice initialization could serve as a warm-up (Zhou et al., 2023) for the training in the new stage, which could accelerate the convergence of this algorithm.

### v2

@@ B.2 Pseudo-code for algorithms
MFC problem (28) (24), $N_{t}$, $N_{z}$, learning rate, weight parameter $\lambda$, number of iterations
@@
              Compute $\mathcal{L}_{\text{HJB}t_{j}}$ through (92)
@@
         compute $z_{t_{j+1}}^{(n)}$, $\tilde{l}_{t_{j+1}}^{(n)}$, $s_{t_{j+1}}^{(n)}$, $H_{t_{j+1}}^{(n)}$ through the forward Euler scheme (9), (11), (12), and (13)
@@
MFC problem (86) (87), $N_{t}$, $N_{z}$, learning rate, weight parameter $\lambda$, number of iterations
@@
              Compute $\mathcal{L}_{\text{HJB}t_{j}}$ through (99)
@@
         compute $z_{t_{j+1}}^{(n)}$ $s_{t_{j+1}}^{(n)}$, $H_{t_{j+1}}^{(n)}$ through the forward Euler scheme (9), (12), and (13)
@@
MFC problem (19) (20), $N_{t}$, $N_{z}$, network structure (8), learning rate, number of iterations, partition of the total time interval $\{I_{m}\}_{m=1}^{N_{\text{stage}}}$
@@ B.3 The errors in the numerical examples
The errors we present in the tables in section 5 are the average errors over $10$ independent runs. Note that there are two sources that contribute to these three errors $\text{err}_{\rho}$, $\text{err}_{f}$, and $\text{err}_{s}$. We take the terminal time $t_{t_{N}}=t_{\text{end}}$ as an example. The first source is the discretization error for the state dynamic, characterized by $z_{t_{\text{end}}}-z_{t_{N_{t}}}$, where $z_{t_{\text{end}}}$ is the end point of the continuous dynamic (3a) and $z_{t_{N_{t}}}$ is its time discretization. The second source comes from the training error of the neural network $f(\cdot,\cdot;\theta)$. When a function is steep, especially the score function $\nabla\log\rho$ in a non-Gaussian example, a small discretization error could be magnified by a large factor due to this large condition number. This is the reason why the score errors for the double well potential function example in Section 5.4 are larger compared with other results.
The errors for the double well potential example. For the double well potential example, it is hard to obtain a reference solution for $t\in(0,t_{\text{end}})$ (see Section 5.4). As an alternative, we compute the error of the velocity field at initial and terminal time $t=0,t_{\text{end}}$ through
@@
where $A(\cdot)$ and $B(\cdot)$ denote the true solutions. Here, we do not present the error of $\theta^{C}$ for two reasons. First, according to (90) and (97), the velocity field does not depend on $\theta^{C}$. Second, we observe from (92) and (99) that a constant shift of $\theta^{C}$ does not affect the regularization loss. Looking more carefully at (91), a constant shift of all the even (or odd) indices of $\theta^{C}$ also does not change the loss (92) or (99). Therefore, it is not meaningful to compute the error for $\theta^{C}$.
@@ B.4 Hyperparameters for the numerical examples
We set the diffusion coefficient $\gamma=1$ for all problems, except for the double well potential example, where $\gamma=0.1$. The reason for this difference is because if we set $\gamma=1$ in the double well example, the terminal distribution $\rho(1,\cdot)$ will be very flat and we will not have a clear bifurcation pattern. In the LQ example in Section 5.3, we set $\beta=0.1$.
@@
For the neural network parametrization (8), we use $\tanh$ as the activation function. The widths of the network are $20$, $50$, and $200$ in $1$, $2$, and $10$ dimensions for the RWPO and LQ problems. For the double moon flow matching problem (in $2$ dimensions), the width is $200$. For the double well example, the widths are $50$, $100$, and $200$ in $1$, $2$, and $10$ dimensions. The double moon flow matching problem and the double well problem are harder due to the bifurcation pattern.
@@
Appendix C remarks
@@
The score dynamic (3c). [11] proposed a similar equation (in equation (13)) before. The higher order equations for the score function are also proposed in [38]. But we have concrete numerical examples. We believe that we are the first to apply the ODE dynamic for the second-order score function (3d) in practice. This equation significantly enhances the accuracy of the regularized MFC solver, as demonstrated in Table 3 and 4.
Parametrization of neural network. In this paper, we parametrize $f$ as a neural network with a single (hidden) layer. As a consequence, the whole map from $z_{0}$ to $z_{t_{\text{end}}}$ is slightly different from the typical structure of a deep neural network, as described in [9]. The layer structure for $f$ is affine-activation-affine, so the composition of multiple layers becomes affine-activation-affine-affine-activation-affine. In contrast, a typical structure for a deep neural network is affine-activation-affine-activation-affine. The composition of two affine functions is still an affine function, so it is not a big difference.
Related work on theoretical MFC. From a theoretical perspective, convergence properties of MFC problems have been extensively studied in [39, 40]. In addition, [41] investigated the limit behavior as the number of agents approaches infinity. The maximum principle, a crucial concept in optimal control, has also been extended to the mean field setting in [42], offering an understanding of the theoretical underpinnings of MFC problems. The mean field game (MFG) problem [43] is closely related to MFC. [12] generalize the HJB equation to the master equation, whose properties have been further investigated in [44, 45]. [46] study the convergence of particle system on Wasserstein space.
HJB regularization for MFC solver. For the HJB regularization technique in B.1, we used absolute value instead of a squared residual in (93) because it performs slightly better. In fact, both forms of regularization could significantly reduce errors.
As for the reason why it reduces errors, we recall that we minimize the discretized loss function in the original Algorithm 1. As a consequence, there is a trend of overfitting to the discretized loss. Adding a regularizer could resolve this overfitting issue. Such issues for discretizations are also discussed in [47].
Computational cost for the score function via (17) and (18). When we compute the score function through (17), the term $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)$ could be tricky. Note that $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)=\nabla_{x}\det(\nabla_{x}T(t,x))/\det(\nabla_{x}T(t,x))$, and the $i$-th component of $\nabla_{x}\det(\nabla_{x}T(t,x))$ can be computed through
@@
Therefore, the $i$-th component of $\nabla_{x}\log\left(\det(\nabla_{x}T(t,x))\right)$ in (17) is
@@
As a consequence, the cost for computing one single score function $s_{t}$ through (17) is either from computing the inverse of Jacobi matrix or solving the related linear system for each dimension. Classical numerical linear algebra methods suggest that the complexity is $\mathcal{O}(d^{3})$, and the cost for one score trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$.
We can also compute the score function through (18), where we assume that the width of the neural network is $\mathcal{O}(d)$, i.e., proportional the dimension. In this way, $\dfrac{\partial z_{t}}{\partial z_{0}}$ and $\dfrac{\partial l_{t}}{\partial z_{0}}$ are computed from auto-differentiations of the deep neural network functions in (15). The computational cost for a single score $s_{t_{j}}$ is $\mathcal{O}(j\,d^{3})$, because we need to compute the back propagation of the deep neural network (15) with depth $j$ using chain rule. However, the total cost for computing a whole trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$ rather than $\mathcal{O}(N_{t}^{2}\,d^{3})$ because we can efficiently store intermediate computations. For example, according to (9), we can compute
@@
with additional $\mathcal{O}(d^{2})$ operations if $\dfrac{\partial z_{t_{j}}}{\partial z_{0}}$ is stored. Similarly, by (10), we can compute
@@
with additional $\mathcal{O}(d^{2})$ operations provided that $\dfrac{\partial l_{t_{j}}}{\partial z_{0}}$ and $\dfrac{\partial z_{t_{j}}}{\partial z_{0}}$ are stored. Note that computing the inverse of the Jacobian matrix or solving the related linear systems requires $\mathcal{O}(d^{3})$ operations following classical numerical linear algebra methods. Therefore, computing $s_{t_{j+1}}$ through (18) requires additional $\mathcal{O}(d^{3})$ operations if we smartly record the results at $t_{j}$. Consequently, the total cost for computing a whole trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$ is $\mathcal{O}(N_{t}\,d^{3})$. We also remark that [48] give a derivation similar to (18).
Other methods to compute the score function. For the KDE method, an approximation for the density function is obtained through taking the kernel convolution with samples directly. Then, one compute its spatial derivative and obtain the score function via $\nabla_{x}\log\rho(t,x)=\nabla_{x}\rho(t,x)/\rho(t,x)$. However, the choice of kernel and bandwidth could be tricky, and the numerical error could be large, especially in high dimensions [49]. Score matching [50] is another commonly used method. With a small trick of integration by part and dropping a constant term that includes the integral with squared score functions, one only requires samples to train a network for the score function. However, training a neural network could be expensive, compared with the KDE method, which evaluates the score function directly.
Regularity for the MFC problem. Formula (3) requires third order derivatives for the velocity field $f$. As a consequence, we need regularity assumption for the MFC problem to guarantee that the ODE system is meaningful. In this work, we assume that $L$ and $F$ are $C^{1,2,2}$ Hölder continuous, and $G$ is $C^{2,2}$ Hölder continuous. According to Schauder estimates [51, Chapter 4], these regularity assumptions guarantee that the velocity field has up to third-order derivatives, ensuring that the ODE system (3) is meaningful.
The multi-stage splicing method. For our multi-stage splicing method (cf. Section 5.2), we partition the total time span evenly and pass the trained neural network from the previous stage into the next stage. When the dynamic becomes complicated, the length of the time interval could also be adapted to suit the problem [52].
Also, the neural network from the previous stage is probably not an accurate velocity for the new stage, but it does contain more information than a random initialization. Therefore, this nice initialization could serve as a warm-start [53] for the training in the new stage, which could accelerate the convergence of this algorithm.